OpenStudy (anonymous):

Multiply. State any excluded values. q+5/-3*-3q/q+2

OpenStudy (anonymous):

is the question \[\frac{ q+5 }{ -3 } * \frac{ -3q }{ q+2 }\]

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

-3 get cancelled \[\frac{ q+5 }{1} * \frac{ q }{ q+2 }\] \[\frac{ q^{2}+5q }{ q+2 }\]

OpenStudy (anonymous):

are you sure because that is not on my answer lists

OpenStudy (anonymous):

so you have \[\frac{q(q+5) }{ q+2 }\]

OpenStudy (anonymous):

My answers are A. -3q^2-15q/-3 B. -3q-15q^2/2q-6 C. q^2-15q/3q D. -3q^2-15q/-3q-6

OpenStudy (anonymous):

D. -3q^2-15q/-3q-6

OpenStudy (anonymous):

i thought you wanted it simplified

OpenStudy (anonymous):

ok ty can you help on my next one im about to post it